# Emerging Technologies for 6G Non-Terrestrial-Networks: From Academia to Industrial Applications

### v1

@@ I INTRODUCTION
Non-terrestrial Network (NTN), encompassing satellites, Unmanned Aerial Vehicles (UAVs), and High-Altitude Platform Stations (HAPS) networks, is a promising solution to address the challenges that terrestrial networks are facing. For instance, through the utilization of satellites and HAPS, NTN can effectively overcome coverage gaps in terrestrial networks, facilitating broadband connectivity across expansive distances. NTN can also play a vital role in enhancing the resilience of terrestrial infrastructure during crises by utilizing UAVs to provide emergency access [5]. Furthermore, by integrating with terrestrial networks, NTN can unlock new possibilities for applications requiring high bandwidth, low latency, and global coverage, including autonomous driving, smart cities, Internet-of-Things (IoT), Augmented Reality, Virtual Reality, cloud gaming, and video conferencing, in the 6G era [6, 7].
TABLE I: Comparison with existing surveys
@@
Despite its potential, the development of NTN is also facing serious challenges. Particularly, NTN often faces significant propagation delay and path loss due to high altitudes. This can significantly hinder the effectiveness of NTN in time-sensitive applications. Moreover, the challenging nature of channel estimation is exacerbated by the time-variant characteristics of NTN. The high moving speed of satellites also causes severe Doppler effects and serious challenges in mobility management. Additionally, the complex integration of NTN and existing terrestrial network infrastructure necessitates careful design and considerations to optimize.
To address these challenges, various solutions have been developed in recent years, leveraging technologies such as network slicing, Artificial Intelligence/Machine Learning (AI/ML), and Open Radio Access Network (ORAN). Particularly, with the powerful ability to create multiple virtual networks from a shared physical network architecture, network slicing enables a single NTN to serve multiple types of applications and users with different demands, e.g., low-latency communication for UAV and high bandwidth for satellite Internet services. Moreover, AI/ML can help NTNs to overcome critical challenges of NTNs. Specifically, AI/ML techniques are very effective in handling non-linear effects and general impairments of channels in NTNs. Those techniques can also enable the autonomous operation of numerous wireless applications, thereby reducing the need for constant human intervention. Additionally, by utilizing the current diverse vendors ecosystem, ORAN can help to significantly improve NTNs’ resilency, scalability, and flexibility.
@@
As summarized in Table I, there are a few surveys on the development of NTN in the literature, such as [6, 7, 4, 9, 10, 8, 11]. Particularly, [8] focuses on the integration of New Radio in NTN, whereas [4] elaborates on the role of NTN in 5G systems. Moreover, [6] surveys the integration of NTN with different types of networks such as IoT, Mobile Edge Computing (MEC), and mmWave. Taking another approach, [7] discusses the utilization of cellular, Wide-Area, and NTN for IoT applications. Additionally, [9] investigates the application of AI/ML in NTN, whereas [10] and [11] discuss the architecture, performance evaluation, and standardization aspects of NTNs and terrestrial networks integration. To the best of our knowledge, there exist no comprehensive studies on how the abovementioned technologies are employed to address the challenges in NTN, especially from an industry perspective. Given the rapid increase in user demands and the emergence of new services, there is an urgent need for new solutions to address the limitations of NTNs. As a result, this paper is expected to fill the gap in the literature and contribute to the future development of 6G networks.
As illustrated in Fig. 1, the rest of this paper is organized as follows. Section II provides a tutorial on NTN and its core enabling technologies including network slicing, ORAN, and AI/ML. Then, the applications of network slicing and AI/ML in NTN are discussed in detail in Section III and IV. Next, the industry applications of ORAN are presented in Section V. Open issues, challenges, and future research directions are presented in Section VI, and conclusions are given in Section VII.
II OVERVIEW OF NTN AND ENABLING TECHNOLOGIES
In this section, we first provide an overview of NTN, including its architectures, advantages, and use cases. Then, the challenges in NTNs are discussed in thoroughly. Finally, we provide a brief background on advanced technologies and infrastructures that can be adopted to efficiently address these challenges, including network slicing, AI/ML, and ORAN.
II-A Overview of NTN
@@
NTNs are communication networks that partially or fully operate through the airborne or spaceborne vehicle(s) [13, 14, 15, 12]. With connections from the air or space, NTNs can empower ultra-reliable communications when the terrestrial infrastructures are not available, such as in remote and unreachable areas or during natural disasters [16]. In addition, NTNs can easily provide multicast connectivity with high data rates and low latency over a large region, enabling massive Machine Type Communication (mMTC) and enhanced Mobile Broadband (eMBB) communications. Moreover, multiple NTNs can be connected together and/or connected to existing terrestrial networks to provide continuous and ubiquitous wireless coverage, and thus bringing us closer to the era of anything, anytime, and anywhere communications [14].
The features, requirements, and protocols for NTNs have been standardized by 3GPP since 2017 with a study item on deployment scenarios and channel models in Release-15 [16]. After that, in Release-16, the required features to enable New Radio (NR) support for NTN are determined by a study in Radio Access Network (RAN) working group. This is then developed further in Release-17 with a set of potential features to enable 5G NR to operate over NTN at frequencies up to 7.125 GHz [16]. Use cases of IoT over NTN with link budget and parameters are also discussed in this Release-17. Release-18 is under development and will focus on 5G systems with satellite backhaul architecture as well as addressing the mobility and continuity problems when connecting different NTNs and also connecting NTNs to terrestrial networks [16].
@@
II-A1 Airborne Platforms
This category includes UAVs and HAPSs. While UAVs operate at low altitudes, (e.g., a few hundred meters), HAPSs, such as airplanes, balloons, and airships, can reach the stratosphere region with altitudes ranging from 20 kilometers to 200 kilometers [12]. One example of airborne NTN is Project Loon [19, 20] from Google which deploys balloons at high altitudes from 18 kilometers to 25 kilometers to provide connectivity to remote and rural areas. Facebook is also involved in airborne NTN by developing solar-powered drones that operate at an altitude of up to 27 kilometers and provide Internet services to an 80-kilometer-radius area below its flying path through Project Aquila [21].
@@ II-A1 Airborne Platforms
II-A2 Spaceborne Platforms
Recently, spaceborne NTN has been emerging as a promising technology for future communication networks (e.g., 5G-Advanced and 6G) where satellites are placed in space to provide communications for users on Earth. These satellites fly around the Earth in specific orbits and can be categorized into: (i) Geostationary Earth Orbit (GEO) and (ii) Non-Geostationary Earth Orbit (NGEO) [12, 22].
@@ II-B Challenges of NTN
  • Channel estimation: Channel estimation is an essential task in every communication system. For NTNs, it is even more challenging due to the inherent time-variant property of NTN platforms. In particular, at high altitudes, NTN platforms fly from horizon to horizon very fast (e.g., around 5-10 minutes for LEO satellites [12]). As a result, users on Earth remain in the coverage of a particular NTN platform for a very short period of time. In addition, the long propagation delay may make estimated channel state information outdated quickly [26]. Hence, traditional estimation methods in terrestrial networks may not be feasible for NTNs, and advanced approaches are required to ensure the success of NTNs.
@@
II-C Technologies and Infrastructures for NTN
Solutions to address the aforementioned challenges of NTNs have been actively developing in the past few years. Among them, network slicing, ORAN, and AI/ML are promising technologies that are expected to play vital roles in NTNs. In this section, we will provide the fundamentals and advantages of these technologies for NTNs.
II-C1 Network Slicing
@@
With the aforementioned advantages, network slicing has been emerging as a potential enabler that can effectively address various challenges of NTNs. A general network slicing-powered NTN system is illustrated in Fig. 3. In particular, like traditional network components, NTN platforms can also be virtualized both in terms of resources and network functions to support different types of services simultaneously. With network slicing, service providers can quickly establish new slices for NTN services, e.g., Earth observation, broadcast services, and broadband services, to serve different groups of users [36]. This can efficiently address the problem of high cost and high delay in building satellite constellations in orbits. In addition, network slicing also allows NTNs to be effectively and flexibly integrated into existing terrestrial networks. Nevertheless, network slicing-powered NTN systems still face several challenges in designs, protocols, and optimizations. These challenges and existing approaches in the literature will be discussed in detail in Section III.
II-C2 ORAN
@@
In Fig. 5, we illustrate the differences in the designs of RAN and ORAN. In particular, ORAN allows the radio access network to work with hardware from any vendor while traditional RAN requires a specific vendor for its hardware. Similarly, in the software part, software from any vendor can be used with a commercial off-the-shelf server, whereas RAN requires proprietary RAN software from a specific vendor. With this open design, ORAN offers various advantages compared to RAN, such as flexibility, open management and orchestration, enabling multi-vendor solutions, and being able to use less-expensive third-party hardware and software. With these advantages, ORAN can be an important technology for NTNs as well as NTN-integrated terrestrial networks. More details applications of ORAN for NTN and current development both in academia and industry will be presented in Section V.
II-C3 AI/ML
AI/ML has been developing significantly since the deep neural network (DNN) architecture was re-invented and trained over a large amount of data by powerful computational computers. It has been successfully applied to various areas such as computer vision, gaming, and natural language processing. In the fields of communications and networking, AI/ML has been emerging as a promising solution to efficiently and significantly improve communication performance. AI/ML is a potential technology for NTNs to overcome their critical challenges that conventional approaches cannot handle well. For instance, deep reinforcement learning, an advanced AI/ML algorithm, can be used to address resource allocation, routing, handover, and beamforming problems in NTNs [40, 41, 42] due to its capability in dealing with the dynamics and uncertainty of the system. Deep learning (DL) with advanced neural network architectures such as Convolutional Neural Network (CNN) and Recurrent Neural Network (RNN) is very effective in handling non-linear effects and general impairments of channels in NTNs. In this section, we will present the fundamentals of common AI/ML techniques that can be used to improve the performance of NTNs. The details of existing AI/ML applications for NTNs will be presented in Section IV.
AI refers to a field of computer science that focuses on building machines (especially computer systems) to perform tasks that require human intelligence such as learning, problem-solving, decision-making, and reasoning. ML is a special subset of AI in which a machine tries to learn a specific task (e.g., image classification, voice recognition, or signal classification and resource allocation in communication networks) and performance metrics such as classification accuracy and performance loss by using only the data collected from the task. ML can be generally categorized into three subsets: (i) supervised learning, (ii) unsupervised learning, and (iii) reinforcement learning [43].
  • Reinforcement learning: This type of ML does not require a prior dataset for training. In particular, reinforcement learning is trained through trial and error by interacting with an external environment. Given a particular state, the reinforcement learning agent makes an action based on its current policy. After that, it observes the next system state and the immediate reward calculated by a predefined reward function. All these observations will be learned by the agent to gradually obtain the optimal policy. The standard techniques in reinforcement learning include Q-learning, deep Q-learning, and multi-armed bandit [43, 44, 45].
Recently, DL, a subset of ML, has been proposed to further improve the prediction performance of ML models by using DNNs. Compared to conventional ML algorithms, DL has several advantages such as no need for system modeling, supporting parallel and distributed algorithms, and being reusable. A typical DNN consists of four main components: (i) neurons, (ii) weights, (iii) biases, and (iv) activation functions. In particular, layers of a DNN are connected to each other by neurons, also known as nodes. Each neuron has an activation function such as tanh, sigmoid, and relu [44]. The activation function is used to calculate the output of each neuron given its weight and bias. During training, the weights of the DNN are updated by calculating the gradient of the loss function. There are three main types of DNNs including (i) Artificial Neural Network (ANN), (ii) Recurrent Neural Network (RNN), and (iii) Convolutional Neural Network (CNN).
Figure 6: General architecture of ANN.
Specifically, ANN, also known as feed-forward neural network, is the most common type of DNNs. Typically, an ANN consists of nonlinear processing layers, including an input layer, several fully connected hidden layers, and an output layer as shown in Fig. 6. As a hidden layer takes the outputs of its previous interconnected layer as its inputs, ANN processes information in one direction from the input layer through the hidden layer to the output layer. Generally, ANN can work well with nonlinear functions, and thus it can be considered a universal function approximator.
Figure 7: General architecture of RNN.
Differently, RNN is a special architecture of DNNs that is widely used for time series data or data that involves sequences. To do that, RNN deploys a feedback loop and hidden states to store information of previous inputs to improve the learning processes of the next data, as illustrated in Fig. 7. In particular, the output of the RNN cell at time $t-1$ will be stored in the hidden state $h_{t}$. This stored information will be used for learning the next sequence at time $t$.
Figure 8: General architecture of CNN.
Finally, CNN is designed mainly for training over image data. The general architecture of a CNN is illustrated in Fig. 8. Specifically, a CNN consists of convolution layers that have a set of convolutional filters. Each convolutional filter extracts specific features from image data. After each convolution operation, CNN uses a Rectified Linear Unit (ReLU) transformation to maintain positive values during training. This helps the training process to be faster and more effective. The pooling layer, also known as downsampling, is used to reduce the number of training parameters by performing dimensionality reduction. It has been widely demonstrated that CNN can handle image data much more effectively than ANN. This is because CNN does not need to convert images to 1-dimensional data before training, which increases the number of training features as well as removes the correlations of features in images. CNN, in contrast, can learn these features directly from image data by using convolutional layers.
III NETWORK SLICING-AIDED NTN

### v2

@@ 1 INTRODUCTION
Non-terrestrial Network (NTN), encompassing satellites, Unmanned Aerial Vehicles (UAVs), and High-Altitude Platform Stations (HAPS) networks, is a promising solution to address the challenges that terrestrial networks are facing. For instance, through the utilization of satellites and HAPS, NTN can effectively overcome coverage gaps in terrestrial networks, facilitating broadband connectivity across expansive distances. NTN can also play a vital role in enhancing the resilience of terrestrial infrastructure during crises by utilizing UAVs to provide emergency access [5]. Furthermore, by integrating with terrestrial networks, NTN can unlock new possibilities for applications requiring high bandwidth, low latency, and global coverage, including autonomous driving, smart cities, Internet-of-Things (IoT), Augmented Reality, Virtual Reality, cloud gaming, and video conferencing, in the 6G era [6, 7]. Table 1 [8] summarizes the main differences between NTNs and terrestrial networks.
Table 1: Comparison between NTNs and Terrestrial Networks [8]
Aspect Non-Terrestrial Networks (NTN) Terrestrial Networks
Coverage Area Global (up to 3500 km) Regional (up to 100km)
Communication Rates Up to 150 Mbps (Starlink) More than 100 Mbps on average
Latency Higher latency (more than 500ms for GEO satellite) Lower latency (lower than 40ms )
Infrastructure Cost Very high for satellites Lower for ground infrastructure
Reliability Robust to disasters Vulnerable to disasters, single points of failure
Accessibility Can reach remote/underserved areas Difficulty in remote area coverage
Mobility Support Seamless mobility across wide areas Frequent handovers for mobility
Table 2: Comparison with existing surveys
@@
Despite its potential, the development of NTN is also facing serious challenges. Particularly, NTN often faces significant propagation delay and path loss due to high altitudes. This can significantly hinder the effectiveness of NTN in time-sensitive applications. Moreover, the challenging nature of channel estimation is exacerbated by the time-variant characteristics of NTN. The high moving speed of satellites also causes severe Doppler effects and serious challenges in mobility management. Another limitation of NTN is the high initial investment cost compared to that of terrestrial networks, especially for satellites, e.g., more than \$2 million launch cost per satellite [14]. Additionally, the complex integration of NTN and existing terrestrial network infrastructure necessitates careful design and considerations to optimize.
To address these challenges, various solutions have been developed in recent years, leveraging technologies such as network slicing, Artificial Intelligence/Machine Learning (AI/ML), and Open Radio Access Network (ORAN). Particularly, with the powerful ability to create multiple virtual networks from a shared physical network architecture, network slicing enables a single NTN to serve multiple types of applications and users with different demands, e.g., low-latency communication for UAVs and high bandwidth for satellite Internet services. Moreover, AI/ML can help NTNs to overcome critical challenges of NTNs. Specifically, AI/ML techniques are very effective in handling non-linear effects and general impairments of channels in NTNs. Those techniques can also enable the autonomous operation of numerous wireless applications, thereby reducing the need for constant human intervention. Additionally, by utilizing the current diverse vendors ecosystem, ORAN can help to significantly improve NTNs’ resiliency, scalability, and flexibility.
@@
As summarized in Table 2, there are a few surveys on the development of NTN in the literature, such as [6, 7, 4, 10, 11, 9, 12]. Particularly, [9] focuses on the integration of New Radio in NTN, whereas [4] elaborates on the role of NTN in 5G systems. Moreover, [6] surveys the integration of NTN with different types of networks such as IoT, Mobile Edge Computing (MEC), and mmWave. Taking another approach, [7] discusses the utilization of cellular, Wide-Area, and NTN for IoT applications. Additionally, although [10] and [13] provide comprehensive surveys on the application of AI/ML in NTN, network slicing and ORAN are not the focus of these surveys. Taking another approach, [11] and [12] discuss the architecture, performance evaluation, and standardization aspects of NTNs and terrestrial networks integration. To the best of our knowledge, there exist no comprehensive studies on how the abovementioned technologies are employed to address the challenges in NTN, especially from an industry perspective. Given the rapid increase in user demands and the emergence of new services, there is an urgent need for new solutions to address the limitations of NTNs. As a result, this paper is expected to fill the gap in the literature and contribute to the future development of 6G networks.
As illustrated in Fig. 1, the rest of this paper is organized as follows. Section 2 provides a tutorial on NTN and its core enabling technologies including network slicing, ORAN, and AI/ML. Then, the applications of network slicing and AI/ML in NTN are discussed in detail in Section 3 and 4. Next, the industry applications of ORAN are presented in Section 5. Open issues, challenges, and future research directions are presented in Section 6, and conclusions are given in Section 7.
2 Overview of NTN and Enabling Technologies
In this section, we first provide an overview of NTN, including its architectures, advantages, and use cases. Then, the challenges in NTNs are discussed thoroughly. Finally, we provide a brief background on advanced technologies and infrastructures that can be adopted to efficiently address these challenges, including network slicing, AI/ML, and ORAN.
2.1 Overview of NTN
@@
NTNs are communication networks that partially or fully operate through the airborne or spaceborne vehicle(s) [16, 8, 17, 15]. With connections from the air or space, NTNs can empower ultra-reliable communications when terrestrial infrastructures are not available, such as in remote and unreachable areas or during natural disasters [18]. In addition, NTNs can easily provide multicast connectivity with high data rates and low latency over a large region, enabling massive Machine Type Communication (mMTC) and enhanced Mobile Broadband (eMBB) communications. Moreover, multiple NTNs can be connected together and/or connected to existing terrestrial networks to provide continuous and ubiquitous wireless coverage, thus bringing us closer to the era of anything, anytime, and anywhere communications [8].
The development of NTNs can be traced back to the 1990s with the commercialization efforts of the two companies Globalstar and Iridium Communications [19]. By deploying several satellites, they can provide low-bandwidth connections to specialized handsets. Since then, NTN technologies have been studied extensively by both academia and industry with a vision of creating a network of satellites and providing Internet services to users anywhere on Earth. The potential features, requirements, and protocols for NTNs have been standardized by 3GPP since 2017 with a study item on deployment scenarios and channel models in Release-15 [18]. After that, the required features to enable New Radio (NR) support for NTN are determined by a study in the Radio Access Network (RAN) working group of Release-16. This is then developed further in Release-17 with a set of potential features to enable 5G NR to operate over NTN at frequencies up to 7.125 GHz [18]. Potential use cases of IoT over NTN with link budget and parameters are also discussed in this Release-17. Release-18 is under development and will focus on 5G systems with satellite backhaul architecture as well as addressing the mobility and continuity problems when connecting different NTNs and also connecting NTNs to terrestrial networks [18].
@@
2.1.1 Airborne Platforms
This category includes UAVs and HAPSs [22]. While UAVs operate at low altitudes, (e.g., a few hundred meters), HAPSs, such as airplanes, balloons, and airships, can reach the stratosphere region with altitudes ranging from 20 kilometers to 200 kilometers [15]. One example of airborne NTN is Project Loon [23, 24] from Google which deploys balloons at high altitudes from 18 kilometers to 25 kilometers to provide connectivity to remote and rural areas. Facebook is also involved in airborne NTN by developing solar-powered drones that operate at an altitude of up to 27 kilometers and provide Internet services to an 80-kilometer-radius area below its flying path through Project Aquila [25].
@@ 2.1.1 Airborne Platforms
2.1.2 Spaceborne Platforms
Recently, spaceborne NTN has been emerging as a promising technology for future communication networks (e.g., 5G-Advanced and 6G) where satellites are placed in space to provide communications for users on Earth. These satellites fly around the Earth in specific orbits and can be categorized into: (i) Geostationary Earth Orbit and (ii) Non-Geostationary Earth Orbit [26, 27, 15, 28].
Geostationary Earth Orbit (GEO)
@@ 2.2 Challenges of NTN
  • Channel estimation: Channel estimation is an essential task in every communication system. For NTNs, it is even more challenging due to the inherent time-variant property of NTN platforms. In particular, at high altitudes, NTN platforms fly from horizon to horizon very fast (e.g., around 5-10 minutes for LEO satellites [15]). As a result, users on Earth remain in the coverage of a particular NTN platform for a very short period. In addition, the long propagation delay may make estimated channel state information outdated quickly [32]. Hence, traditional estimation methods in terrestrial networks may not be feasible for NTNs, and advanced approaches are required to ensure the success of NTNs.
@@
2.3 Technologies and Infrastructures for NTN
Solutions to address the aforementioned challenges of NTNs have been actively developing in the past few years, including beamforming designs, efficient resource allocation, dynamic routing, and intelligent operation management [34, 35], by leveraging advanced technologies in AI/ML, communications and networking, and computing. Among them, network slicing, ORAN, and AI/ML are promising technologies that are expected to play vital roles in NTNs. In this section, we will provide the fundamentals and advantages of these technologies for NTNs.
2.3.1 Network Slicing
@@
With the aforementioned advantages, network slicing has emerged as a potential enabler that can effectively address various challenges of NTNs. A general network slicing-powered NTN system is illustrated in Fig. 3. In particular, like traditional network components, NTN platforms can also be virtualized both in terms of resources and network functions by using SDN and NFV to support different types of services simultaneously. With network slicing, service providers can quickly establish new slices for NTN services such as Earth observation, broadcast services, and broadband services to serve different groups of users [44]. This can efficiently address the problem of high cost and high delay in building satellite constellations in orbits. In addition, network slicing also allows NTNs to be effectively and flexibly integrated into existing terrestrial networks. Nevertheless, network slicing-powered NTN systems still face several challenges in designs, protocols, and optimizations. In particular, the high mobility of satellites can create frequent topology changes and handovers that may not be fully addressed by using existing network slicing technologies. In addition, satellite communications have higher latency and energy consumption compared to traditional communications. As such, managing and allocating different types of virtual resources to maintain efficient network slicing operations are more complex in NTNs than in terrestrial networks. These challenges and existing approaches in the literature will be discussed in detail in Section 3.
2.3.2 AI/ML
AI/ML has been developing significantly since the deep neural network (DNN) architecture was re-invented and trained over a large amount of data by powerful computational computers. It has been successfully applied to various areas such as computer vision, gaming, and natural language processing. In the fields of communications and networking, AI/ML has been emerging as a promising solution to efficiently and significantly improve communication performance. AI/ML is a potential technology for NTNs to overcome their critical challenges that conventional approaches cannot handle well. For instance, Deep Reinforcement Learning (DRL), an advanced AI/ML algorithm, can be used to address resource allocation, routing, handover, and beamforming problems in NTNs [45, 46, 47] due to its capability in dealing with the dynamics and uncertainty of the system. Deep learning (DL) with advanced neural network architectures such as Convolutional Neural Network (CNN) and Recurrent Neural Network (RNN) is very effective in handling non-linear effects and general impairments of channels in NTNs. In this section, we will present the fundamentals of common AI/ML techniques that can be used to improve the performance of NTNs. The details of existing AI/ML applications for NTNs will be presented in Section 4.
AI refers to a field of computer science that focuses on building machines (especially computer systems) to perform tasks that require human intelligence such as learning, problem-solving, decision-making, and reasoning. ML is a special subset of AI in which a machine tries to learn a specific task (e.g., image classification, voice recognition, or signal classification and resource allocation in communication networks) and performance metrics such as classification accuracy and performance loss by using only the data collected from the task. ML can be generally categorized into three subsets: (i) supervised learning, (ii) unsupervised learning, and (iii) Reinforcement Learning (RL) [48].
  • Reinforcement learning: This type of ML does not require a prior dataset for training. In particular, RL is trained through trial and error by interacting with an external environment. Given a particular state, the RL agent makes an action based on its current policy. After that, it observes the next system state and the immediate reward calculated by a predefined reward function. All these observations will be learned by the agent to gradually obtain the optimal policy. The standard techniques in RL include Q-learning, deep Q-learning, and multi-armed bandit [48, 49, 50, 51].
To further improve the prediction performance of ML models, DL has been proposed by leveraging the capabilities of DNNs in learning from a large amount of data. Compared to conventional ML algorithms, DL has several advantages such as no need for system modeling, supporting parallel and distributed algorithms, and being reusable. A typical DNN consists of four main components: (i) neurons, (ii) weights, (iii) biases, and (iv) activation functions. In particular, layers of a DNN are connected to each other by neurons, also known as nodes. Each neuron has an activation function such as tanh, sigmoid, and relu [49]. The activation function is used to calculate the output of each neuron given its weight and bias. During training, the weights of the DNN are updated by calculating the gradient of the loss function. There are three main types of DNNs including (i) Artificial Neural Network (ANN), (ii) Recurrent Neural Network (RNN), and (iii) Convolutional Neural Network (CNN).
  • Artificial Neural Network: ANN, also known as feed-forward neural network, is the most common type of DNNs. Typically, an ANN consists of nonlinear processing layers, including an input layer, several fully connected hidden layers, and an output layer as shown in Fig. 4. As a hidden layer takes the outputs of its previous interconnected layer as its inputs, ANN processes information in one direction from the input layer through the hidden layer to the output layer. Generally, ANN can work well with nonlinear functions, and thus it can be considered a universal function approximation. Due to its simple architecture and ability to extract useful information from training data, ANN has been commonly adopted to address emerging issues in communications and networking. For example, the authors in [52] propose to use an ANN architecture with only three hidden layers for channel estimation and signal detection in orthogonal frequency-division multiplexing (OFDM) systems. Simulation and experimental results then reveal that ANN is a promising tool for channel estimation and signal detection in wireless environments under complex channel distortion and interference. Moreover, in [53], the authors present several use cases of applying ANNs to different problems in wireless communications such as UAV-based wireless networks, radio access, and mobile edge caching and computing.
    Figure 4: General architecture of ANN.
  • Recurrent Neural Network: RNN is a special architecture of DNNs that is widely used for time series data or data that involves sequences. To do that, RNN deploys a feedback loop and hidden states to store information of previous inputs to improve the learning processes of the next data, as illustrated in Fig. 5. In particular, the output of the RNN cell at time $t-1$ will be stored in the hidden state $h_{t}$. This stored information will be used for learning the next sequence at time $t$. In complex systems, RNN may not perform well due to the “vanishing” or “exploding” gradient problem during the backpropagation operation. To overcome this issue, an extended version of RNN, namely Long Short-Term Memory (LSTM), is proposed. In particular, LSTM employs additional gates to determine the amount of previous information in the hidden state that will be used for the output and the next hidden state. In this way, LSTM can efficiently learn the long-term dependencies in training data while mitigating the issue of the backpropagation process. RNN, especially LSTM, has emerged as a promising architecture for signal classification in wireless communications due to the fact that signals are naturally sequential and collected over multiple antennas [54, 55]. In addition, LSTM can be used for resource allocation, modulation classification, intrusion detection, and beamforming [56, 57, 58].
    Figure 5: General architecture of RNN.
  • Convolutional Neural Network: CNN is designed mainly for training over image data. The general architecture of a CNN is illustrated in Fig. 6. Specifically, a CNN consists of convolution layers that have a set of convolutional filters. Each convolutional filter extracts specific features from image data. After each convolution operation, CNN uses a Rectified Linear Unit (ReLU) transformation to maintain positive values during training. This helps the training process to be faster and more effective. The pooling layer, also known as downsampling, is used to reduce the number of training parameters by performing dimensionality reduction. It has been widely demonstrated that CNN can handle image data much more effectively than ANN. This is because CNN does not need to convert images to 1-dimensional data before training, which increases the number of training features as well as removes the correlations of features in images. CNN, in contrast, can learn these features directly from image data by using convolutional layers. As a result, CNN has been widely applied in communications and networking to handle data in the form of images or high-dimensional matrices. For example, the authors in [59] propose to use CNN for spectrum sensing in cognitive radio networks. In addition, CNN can be used for automatic modulation classification as demonstrated in [60] by taking IQ time-domain vectors of modulated signals as its inputs. Due to the ability to handle high-dimensional input data, CNN is a promising architecture for channel feature extraction as studied in [61].
    Figure 6: General architecture of CNN.
As mentioned, DL approaches usually require sufficient training data from users to achieve good performance. However, in practice, users may not be willing to share their data with the centralized server due to privacy and security concerns. To address this issue, federated learning (FL) was introduced in [62]. In FL, each user, i.e., client, uses their local data to train a DL model. After that, they send their model updates, i.e., their models’ weights, to a centralized server for aggregation. The server then sends the aggregated global model to all the clients. Based on this global model, the clients then continue to train it with their local data. This process is repeated until a desirable accuracy is obtained. In this way, FL can effectively address the privacy problem as well as reduce the required bandwidth as only models’ weight is transmitted to the server instead of raw training data. Due to these advantages, FL has been widely adopted in wireless communications and networking to address a wide range of problems such as spectrum management, caching, and IoT [63, 64]. Applications of FL in NTNs will be also discussed in detail in Section 4.
2.3.3 ORAN
@@
In Fig. 8, we illustrate the differences in the designs of RAN and ORAN. In particular, ORAN allows the radio access network to work with hardware from any vendor while traditional RAN requires a specific vendor for its hardware. Similarly, in the software part, software from any vendor can be used with a commercial off-the-shelf server, whereas RAN requires proprietary RAN software from a specific vendor. With this open design, ORAN offers various advantages compared to RAN, such as flexibility, open management, and orchestration, enabling multi-vendor solutions, and being able to use less-expensive third-party hardware and software. With these advantages, ORAN can be an important technology for NTNs as well as NTN-integrated terrestrial networks. More detailed applications of ORAN for NTN and current development both in academia and industry will be presented in Section 5.
3 Network Slicing-aided NTN